When is marginal cost at its minimum
1
Expert's answer
2019-11-11T15:25:49-0500

Marginal Cost (MC) is the amount added to the Total Cost while producing the extra one unit."MC= \frac{Difference in total cost} {Difference in production}"

"MC=\frac {\triangle{TC}} {\triangle{Q}}"

MC is slope of Total Cost (TC) curve.


Also "MC= \frac{Difference in TVC} {Difference in production}"

"MC=\frac {\triangle{TVC}} {\triangle{Q}}"




In the short run after AVC curve reach its minimum point ATC curve reach its minimum point. Because ATC includes both AVC and AFC. MC curve starts from the below of ATC and AVC curves and when ATC and ATC raising the MC curve raising above the ATC and AVC curves through crossing the minimum points of those curves. MC is equal to the AVC when MC is at its minimum level.

MC is U shape curve as a firm initially starts to increases it's outputs, TC and TVC at a diminishing rate. Due to economies of scale and diminishing returns, MC is decreasing until it reach the minimum point. when firm increasing the production the MC will increase. so do the MC curve.